# Evolution of LISA Observables for Binary Black Holes Lensed by an SMBH

### v1

@@ 2.2 Redshift and Magnification Interpolation
Figure 3: Contour plots showing the redshift and magnification sampled across the disk ranging from $12R_{g}$ to $120R_{g}$ for two separate inclination angles of $65^{\circ}$ and $88^{\circ}$. The observer is looking at the disk from an azimuthal angle of $\theta=0^{\circ}$ and the inclination angle is defined such that the $\varphi=0^{\circ}$ case is face on. At higher inclination angles, the top of the disk tilts away relative to the observer’s perspective. Note the difference in scales between the $65^{\circ}$ and $88^{\circ}$ cases for both redshift and magnification.
@@
Figure 4: Plots showing the interpolated change in redshift and magnification vs azimuthal angle in the disk at gravitational radii of $20R_{g}$ and $100R_{g}$ for separate inclination angles of $65^{\circ}$ and $88^{\circ}$. Note the difference in maxima for the magnification between the two cases.
@@ 2.3 Binary Evolution
The orbital period $P$ for a binary orbiting around an SMBH at radius $r$ in units of $R_{g}$ is given by $P(r)=2\pi GM_{\rm{SMBH}}r^{3/2}c^{-3}$. The azimuth at which we sample our interpolated redshift and magnification can be found as $\theta=(2\pi TP^{-1})\bmod{2\pi}$ where $T$ is the total elapsed time.
We utilize LEGWORK, an open source project developed to predict the evolution of binary sources due to gravitational wave emission (Wagg et al., 2022a, b), producing a baseline for the evolving observed frequency, strain, and characteristic strain. Using the strain amplitude $h_{0}$, as given by LEGWORK for a theoretical observer at a distance of $410$  Mpc, we find the magnified strain amplitude as $h_{\mu}=h_{0}\sqrt{\mu}$ where $\mu$ is the magnification factor given by the position of the binary along its orbit.
@@ 3 Results
Figure 5: Characteristic Strain vs Frequency for a binary at $20R_{g}$ and $100R_{g}$ in disk with inclinations of $65^{\circ}$ and $88^{\circ}$. The two peaks correspond to the where the rate of change of the redshift is small. The minima of the evolution corresponds to when the binary is positioned directly in front of or behind the SMBH, where magnification is at a minimum or maximum, respectively. The $20R_{g}$ orbit would have a period of $\approx\text{3 days}$, with the $100R_{g}$ orbit having a period of $\approx\text{35 days}$. This would correspond to a burst-like signal occurring every 1.5 and 17.5 days respectively.
Using the interpolated redshift and magnification, shown in Figure 4, to approximate the characteristic strain throughout the evolution of the binary allows us to plot the characteristic strain vs frequency over the LISA sensitivity curve, as seen in Figure 5. The evolution of the isolated binary in a vacuum and assuming pure GR evolution, can be seen in magenta. The first orbit of the binary around the SMBH can be seen in orange, with the entire lifespan of the binary in the LISA frequency range seen in blue.
@@
The dual peaks observed in Figure 5, therefore correspond to portions of the orbit where there is some magnification and the change in redshift/frequency is slowest ($\dot{f}_{\rm obs}$ is near zero). This can be seen around an azimuth of $\varphi\approx\pi/2$ and $\varphi\approx 3\pi/2$ in Fig. 4.
We note several additional important trends visible in Fig. 5: going from $r=20$ to 100 $R_{g}$ we see the characteristic strains are larger at the same frequencies since the change in $\dot{f}_{\rm{obs}}$ is slower as a result of a larger, slower orbit. Going from an inclination of $\theta=65$ to $88$ degrees, the trace of the plot during an orbit is different since the magnification factor is higher when the binary is behind the SMBH in the more inclined disk. This feature is present in both $r=20$ and 100 $R_{g}$ orbits. Finally, in all cases we note the sharp burst-like quality of the observed lensed signals in the higher frequency region of the LISA band, where background sources are expected to be less common. We will return to this latter point in the discussion section.
@@ 4.1 Rates and LISA Detectability
In preparation for when LISA starts observing, we want to build a more complete picture of our universe so that we can either confirm the accuracy existing models, or know that new theories need to be developed. To start in this endeavor, we can look at existing rate calculations for the LVK population. The occurrence of BBH mergers near the last migration trap is analytically estimated to be $\sim 0.4~{}\text{Gpc}^{-3}\text{yr}^{-1}$ for a $3\times 10^{8}M_{\odot}$ SMBH in an AGN with a lifespan of $10^{7}~{}\text{yrs}$ (Peng & Chen, 2021). One way to both test and extend this estimate for LISA would be to utilize a population synthesis code such as McFACTS (McKernan et al., 2024). Depending on the properties of the AGN disk used to run McFACTS, one would be able to count the occurrence of objects fitting our criteria and assume a rate at which we expect LISA to observe these types of objects.
@@ 4.3 Intrinsic Properties of the Disk
Detecting a binary as we described, in an AGN disk, could provide an extremely useful probe of the underlying disk morphology. In our simplified case, it is apparent that one would be able to infer the inclination of the disk (at the orbit of the binary) based on the change in strain amplitude between the binary passing in front of and behind the SMBH. A larger $\Delta$strain would be produced by a more edge-on orientation (as is apparent by comparing the lower panels to the upper panels in Fig. 5. On the other hand, a precisely face-on orientation would produce no magnification or time-varying Doppler shift (though there would still be a constant frequency offset from the emitted to observed frequency, due to gravitational redshifting). If the binary is sufficiently close to the SMBH, we would expect the emitted gravitational waves incur effects from relativistic beaming. In such a case, we would expect to see a difference between the local maxima of the two peaks of the characteristic strain, therefore one could determine the phase alignment between the GW observables and other possible observables of the system.
Support was provided by Schmidt Sciences, LLC. for JP, HB & MO. KESF & BM are supported by NSF AST- 2206096, NSF AST- 1831415 and Simons Foundation Grant 533845 as well as Simons Foundation sabbatical support. The Flatiron Institute is supported by the Simons Foundation. KESF wishes to acknowledge extremely helpful conversations with Xian Chen and Laura Sberna at New Ideas on the Origin of Black Hole Mergers at NBIA, in August 2024. HB acknowledges the support from the Czech Science Foundation (GACR) Junior Star grant no. GM24-10599M.

### v2

@@ 2.2 Redshift and Magnification Interpolation
Figure 3: Contour plots showing the redshift and magnification sampled across the disk ranging from $12R_{g}$ to $120R_{g}$ for two separate inclination angles of $65^{\circ}$ and $88^{\circ}$. The observer is looking at the disk from an azimuthal angle of $\theta=0^{\circ}$ and the inclination angle is defined such that the $\varphi=0^{\circ}$ case is face-on. At higher inclination angles, the top of the disk tilts away relative to the observer’s perspective. Note the difference in scales between the $65^{\circ}$ and $88^{\circ}$ cases for both redshift and magnification.
@@
Figure 4: Plots showing the interpolated change in redshift and magnification vs. azimuthal angle in the disk at gravitational radii of $20R_{g}$ and $100R_{g}$ for separate inclination angles of $65^{\circ}$ and $88^{\circ}$. Note the difference in maxima for the magnification between the two cases.
@@ 2.3 Binary Evolution
The orbital period $P$ for a binary orbiting around an SMBH at radius $r$ in units of $R_{g}$ is given by $P(r)=2\pi GM_{\rm{SMBH}}r^{3/2}c^{-3}$. The azimuth at which we sample our interpolated redshift and magnification can be found as $\varphi=(2\pi TP^{-1})\bmod{2\pi}$ where $T$ is the total elapsed time.
We utilize LEGWORK, an open source project developed to predict the evolution of binary sources due to gravitational wave emission (Wagg et al., 2022a, b), producing a baseline for the evolving observed frequency, strain, and characteristic strain. To find the magnified strain amplitude $h_{\mu}$, the strain amplitude $h_{0}$ is multiplied by the square of the magnification factor $\sqrt{\mu}$, given by the position of the binary along its orbit. The strain amplitude $h_{0}$ is obtained via LEGWORK for a theoretical observer at a distance of 410410  Mpc.
@@ 3 Results
Figure 5: Characteristic Strain vs. Frequency for a binary at $20R_{g}$ and $100R_{g}$ in disk with inclinations of $65^{\circ}$ and $88^{\circ}$. The two peaks correspond to the where the rate of change of the redshift is small. The minima of the evolution corresponds to when the binary is positioned directly in front of or behind the SMBH, where magnification is at a minimum or maximum, respectively. The $20R_{g}$ orbit would have a period of $\approx\text{3 days}$, with the $100R_{g}$ orbit having a period of $\approx\text{35 days}$. This would correspond to a burst-like signal occurring every 1.5 and 17.5 days respectively.
Using the interpolated redshift and magnification, shown in Figure 4, to approximate the characteristic strain throughout the evolution of the binary allows us to plot the characteristic strain vs. frequency over the LISA sensitivity curve, as seen in Figure 5. The evolution of the isolated binary in a vacuum and assuming pure GR evolution, can be seen in magenta. The first orbit of the binary around the SMBH can be seen in orange, with the entire lifespan of the binary in the LISA frequency range seen in blue.
@@
The dual peaks observed in Figure 5, therefore correspond to portions of the orbit where there is some magnification and the change in redshift/frequency is slowest ($\dot{f}_{\rm obs}$ is near zero). This can be seen around an azimuth of $\varphi\approx\pi/2$ and $\varphi\approx 3\pi/2$ in Figure 4.
We note several additional important trends visible in Figure 5: going from $r=20$ to 100 $R_{g}$ we see the characteristic strains are larger at the same frequencies since the change in $\dot{f}_{\rm{obs}}$ is slower as a result of a larger, slower orbit. Going from an inclination of $\theta=65$ to $88$ degrees, the trace of the plot during an orbit is different since the magnification factor is higher when the binary is behind the SMBH in the more inclined disk. This feature is present in both $r=20$ and 100 $R_{g}$ orbits. Finally, in all cases we note the sharp burst-like quality of the observed lensed signals in the higher frequency region of the LISA band, where background sources are expected to be less common. We will return to this latter point in the discussion section.
@@ 4.1 Rates and LISA Detectability
In preparation for when LISA starts observing, we want to build a more complete picture of our universe so that we can either confirm the accuracy existing models, or know that new theories need to be developed. To start in this endeavor, we can look at existing rate calculations for the LVK population.
The occurrence of BBH mergers near the last migration trap is analytically estimated to be
$R=\frac{n_{\rm QSO}}{2}\frac{N_{\rm BH}}{\tau_{\rm AGN}}\sim 0.4~\text{Gpc}^{-3}\text{yr}^{-1}$ (2)
where $n_{\rm QSO}=10^{-7}~\text{Mpc}^{-3}$ is the number density of QSOs, $N_{\rm BH}=6\times 10^{4}$ is the number of black holes in the NSC, and $\tau_{\rm AGN}=10^{7}~\text{yrs}$ is the assumed AGN lifetime (Peng & Chen, 2021). However, this estimate implicitly assumes an aggressive prescription for the disk capture rate of BH from the NSC—nearly all of the BH are captured by the end of the AGN phase. On the other hand, it also assumes a conservative number density of QSO (only the most luminous) which does not account for the full spectrum of BBH forming QSOs and AGN (excluding lower mass & luminosity) that could exist in a detection volume out to $z\simeq 1$. These choices work out such that assuming the opposite (a less aggressive capture rate, but high number density of AGN) could yield a similar answer; additionally the AGN lifetime is sufficiently uncertain that it contributes to uncertainty in our rate estimate and we could imagine a rate that varies by an order of magnitude (or more) in either direction.
One way to both test and extend this estimate for LISA would be to utilize a population synthesis code such as McFACTS (McKernan et al., 2024). Depending on the properties of the AGN disk used to run McFACTS, one would be able to count the occurrence of objects fitting our criteria and assume a rate at which we expect LISA to observe these types of objects. McFACTS currently does not permit binary formation in the inner disk (even well outside the last migration trap), however we can determine a rough estimate for the rate by noting that in a typical run, $\sim 2\%$ of all BH end up in the inner disk and could form binaries at the last migration trap (e.g. McKernan et al., 2024). Assuming all BBH mergers originate from the AGN channel, we would expect at most $2\%$ of the LVK merger rate could come from these strongly lensed binaries or $\sim 0.5~\text{Gpc}^{-3}\text{yr}^{-1}$. This is notably similar to the rate from Peng & Chen (2021), but as with their estimate, there remain at least order of magnitude uncertainties.
@@ 4.3 Intrinsic Properties of the Disk
Detecting a binary as we described, in an AGN disk, could provide an extremely useful probe of the underlying disk morphology. In our simplified case, it is apparent that one would be able to infer the inclination of the disk (at the orbit of the binary) based on the change in strain amplitude between the binary passing in front of and behind the SMBH. A larger change in strain would be produced by a more edge-on orientation (as is apparent by comparing the lower panels to the upper panels in Fig. 5. On the other hand, a precisely face-on orientation would produce no magnification or time-varying Doppler shift (though there would still be a constant frequency offset from the emitted to observed frequency, due to gravitational redshifting). If the binary is sufficiently close to the SMBH, we would expect the emitted gravitational waves incur effects from relativistic beaming. In such a case, we would expect to see a difference between the local maxima of the two peaks of the characteristic strain, therefore one could determine the phase alignment between the GW observables and other possible observables of the system.
Support was provided by Schmidt Sciences, LLC. for JP, HB & MO. KESF & BM are supported by NSF AST- 2206096, NSF AST- 1831415 and Simons Foundation Grant 533845 as well as Simons Foundation sabbatical support. The Flatiron Institute is supported by the Simons Foundation. HB acknowledges the support from the Czech Science Foundation (GACR) Junior Star grant no. GM24-10599M. KESF wishes to acknowledge extremely helpful conversations with Xian Chen and Laura Sberna at New Ideas on the Origin of Black Hole Mergers at NBIA, in August 2024. JP, KESF, and BM wish to acknowledge the useful conversations and abundance of help provided by Tom Wagg throughout this exploration. The authors also wish to acknowledge the helpful suggestions and advice provided by the referee in their evaluation of the manuscript of this article.